OpenStudy (anonymous):

Cot x Sec^4 x = Cot x + 2tan x + tan^3 x

OpenStudy (mertsj):

Is this an identity to be proven?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Does your teacher care which side you work on?

OpenStudy (anonymous):

They said to have me match the right side with the left, but I don't think it matters.

OpenStudy (mertsj):

have you tried anything yet?

OpenStudy (anonymous):

I was thinking I might be able to factor it.

OpenStudy (anonymous):

to me, I replace sec^4 x = (1 = tan^2)^2 and then expand it, and then time cot , and get the RHS

OpenStudy (anonymous):

1+tan^2 not 1=tan^2 sorry

OpenStudy (anonymous):

thank you so much! that helps!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i don't actually see how that works out. i get an extra 1 and the cotx is still multiplying by what is inside the factored out equation.

OpenStudy (mertsj):

\[\cot x(1+\tan ^2x)^2=\cot x(1+2\tan ^2x+\tan ^4x)=\cot x \cot +2\cot x \tan ^2x+\cot x \tan ^4x\]

OpenStudy (mertsj):

Hang on...I'll fix that

OpenStudy (mertsj):

\[\cot x+2\cot x \tan ^2x+\cot x \tan ^4x\]

OpenStudy (mertsj):

Do you agree so far?

OpenStudy (mertsj):

Hoa...any mistakes there?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So we have the cotx. Let's see if we can show that the other two terms are equal to the two terms on the right.

OpenStudy (mertsj):

\[2\cot x \tan ^2x=2\cot x \tan x \tan x=2\tan x\]

OpenStudy (mertsj):

Do you see that?

OpenStudy (mertsj):

Now you do the same thing to the last term

OpenStudy (mertsj):

And you will be done.

OpenStudy (anonymous):

thank you! I will do that!

OpenStudy (mertsj):

you're welcome and thanks to Hoa as well

OpenStudy (anonymous):

thank you both!